OpenStudy (anonymous):

f(x)=1/4x+7. Find x so that f(x)=9

OpenStudy (anonymous):

your job is to solve \[\frac{1}{4x+7}=9\] start with \[1=9(4x+7)\] then it should be straightforward

OpenStudy (precal):

follow satellite73's instructions, they are correct

OpenStudy (anonymous):

@satellite73 so what's the final answer?

OpenStudy (precal):

you tell us and we will verify it for you